Main INDEX
Monthly INDEX
PREV
NEXT
Make New Entry,
Make Followup Entry
User name V. Sulkosky
Log entry time 18:45:49 on February 26, 2009
Entry number 263544
This entry is a followup to: 263312
Followups:
keyword=Experiment Summary for Feb 25, Owl - Feb 25, Swing
Summary for Feb 25 owl - Feb 25 swing
Smooth data taking at fourth pass.
Beam in Use
shift |
ABU |
BANU |
BNA |
2009/02/25 owl | 7.34 |
0.06 | 0.61 |
2009/02/25 day | 5.10 |
2.25 | 0.64 |
2009/02/25 swing | 7.36 |
0.07 | 0.55 |
total | 19.8 |
2.4 | 1.8 |
Main events/issues:
- February 25th
- Production on kinematic 4J: Plhrs = 1.51 GeV/c,
45o.
- VESDA alert getting worse, four indicators blinking red at the
alert level.
- 09:00-11:20 Controlled access to investigate cause for VESDA
alert.
- Compressor turned off during access.
- By end of day shift, alert levels improved significantly.
- 11:20 Resume production on kinematic 4J.
- Continue production on kinematic 4J to the end of swing shift.
Plans for the next couple of days:
- Production for kinematic 4J and 4C until 9 am on February
26th.
- February 26th from 09:00 to 15:00
- Beam studies
- Controlled access for spectrometer magnet work
- Pass change to 5th, wien angle to 30o
- Resume polarized production at 5th pass.
A copy of this log entry has been emailed to: brads,meziani,jpchen,choi,jiang